Give examples if economic problems that are solved with the help of first and second order difference equation.


1
Expert's answer
2021-04-13T07:15:00-0400

i) Utility max problem;"U=Q1Q2"

Subject to; "10Q1+2Q2=240"

Lagrange function

"Z = Q1Q2+ \u03bb(240-10Q1-2Q2)"

b) first-order conditions:

"ZQ1 = Q2− λ10 = 0\\


 ZQ2 = Q1− λ 2 = 0 \\


Zλ = 240 − 10Q1 −2 Q2 =0."


"Z\lambda=240-10Q1-2Q2=0"

"ZQ1=Q2- \lambda10=0"

"ZQ2=Q1-\lambda2=0"


"\begin{bmatrix}
 0 & -10 & -2 \\
 -10 & 0 &1 \\
-2 & 1 & 0
\end{bmatrix}" "\begin{bmatrix}
 \lambda \\
 Q1 \\ Q2
\end{bmatrix}" = "\begin{bmatrix}
 -240 \\
 0 \\ 0
\end{bmatrix}"


Q1M"=\frac{240}{2[-10]}= -12"


Q2M"=\frac{240}{2[-2]}=-60"


"\\lambda=\\frac{240}{2[-10.-2] }\n\u200b\t\n =6"


ii.)Second order condition.